OpenStudy (anonymous):

If there is a 30% chance of sun tomorrow and a 40% chance of wind and no sun, what is the probability that it is windy, given that it is not sunny? Round your answer to the nearest percent

OpenStudy (kropot72):

\[\large P(S)=0.3\] \[\large P(\bar{S})=1-0.3=0.7\] \[\large P(W \cap \bar{S})=0.4\] \[\large P(W|\bar{S})=\frac{P(W\cap\bar{S}}{P(\bar{S})}=\frac{0.4}{0.7}=\ you\ can\ calculate\] When you have calculated the probability, multiply the result by 100 to convert to a percentage and round to the nearest percent.